OpenStudy (anonymous):

2x/7+4/5=5 read as 2x over 7 + 4 over 5 equals 5

OpenStudy (anonymous):

\[\frac{ 2x }{ 7 } + \frac{ 4 }{ 5 } = 5\] \[\frac{ 2x }{ 7 } + \frac{ 4 }{ 5 } = \frac{ 25 }{ 5 }\] \[\frac{ 2x }{ 7 } = \frac{ 21 }{ 5 }\] \[2x = \frac{ 147 }{ 5 }\] \[x = \frac{ 147 }{ 10 }\] :)

OpenStudy (anonymous):

Hope this helps!

OpenStudy (anonymous):

you are such a great help !!!!!!!! i could just hug you right now!